Brenan, Inc. purchased equipment at the beginning of 2004 for $2,100,000. Brenan. The equipment has an estimated residual value (salvage value) of $100,000 and an estimated life of 5 years or 100,000 hours of operation. The machinery was operated for 15,000 hours in 2004, 20,000 hours in 2005, 35,000 hours in 2006, 20,000 hours in 2007, and 10,000 hours in 2008 Record the depreciation journal entry using the DOUBLE-DECLINING METHOD for 2006
Answer #1

Solution

Brenan Inc

Journal entry to record depreciation expense under double-declining balance method for 2006:

Date

Account Titles and Explanation

Debit

Credit

December 31, 2006

Depreciation Expense - Equipment

$302,400

Accumulated Depreciation - Equipment

$302,400

(To record depreciation expense for 2006)

                                         

Computations:

For double declining balance method,

Depreciation expense = 200% straight line depreciation rate x cost

Straight line depreciation rate = 1/useful life

Useful life of equipment = 5 years

Straight line depreciation rate = 1/5 = 20%

DDB depreciation rate = 2 x 20% = 40%

Cost of equipment = $2,100,000

Date of purchase = beginning of 2004

Depreciation expense for 2004 –

= 2,100,000 x 40% = $840,000

Book value at end of year = cost – accumulated depreciation

Book value at end of 2004 = 2,100,000 – 840,000 = $1,260,000

Depreciation expense, 2005 = book value x depreciation rate

= 1,260,000 x 40% = $504,000

Accumulated depreciation at end of 2005 = 840,000 + 504,000 = $1,344,000

Book value at end of 2005 = 2,100,000 – 1,344,000 = $756,000

Depreciation expense 2006 = 756,000 x 40% = $302,400

Accumulated depreciation, 2006 = 840,000 + 504,000 + 302,400 = $1,646,400

Book value at end of year 2006 = 2,100,000 – 1,646,400 = $453,600